# Bayesian Transfer Operators in Reproducing Kernel Hilbert Spaces

### v1

@@ 3.1 Generalization Errors
3.2 Decoupling the Lifted Observation Model
To assess the sensitivity of our algorithm to the approximation introduced in Definition 2.1, we separately train a regularization parameter for the lifted model using the kernelized targets $\mathbf{K}_{ZY}$ (after estimating the optimal kernel hyperparameters and pseudo-inputs as discussed in Section 2.5). In effect, we are defining the RKHS noise model with $\sigma_{\mathcal{Y}}^{2}\kappa_{\textrm{bc}}\left(\cdot,\cdot\right)$ where $\sigma_{\mathcal{Y}}^{2}$ is not directly coupled to the variance of the sensor noise in the state space, $\sigma_{Y}^{2}$. Put differently, the regularization term in (2.5) corresponds to $\sigma_{\mathcal{Y}}$, whereas in (2.7) it corresponds to $\sigma_{Y}$. This establishes a division of roles between the two regularization terms: $\sigma_{Y}$ is predominantly responsible for projecting to the state space, while $\sigma_{\mathcal{Y}}$ controls the accuracies of the multi-step forecasts. Consequently, the relationship in (2.11) does not hold anymore, but our model is more flexible.
The effect of decoupling is illustrated in Figure 3 where we observe that the growth rate of the GP-TCCA model has markedly reduced, leading to more accurate long-term predictions. We therefore increase the credence in our hypothesis: rigorously accounting for the distortion introduced by the nonlinear feature map will yield a more robust model. For example, we are naturally tempted to use an input-dependent heteroskedastic covariance structure for $\nu_{i}$. This is apparent when we take into consideration that $\kappa(\bm{x},Y)$ is typically a nonlinear many-to-one transformation of $Y$.
@@ 3.3 Reprojections
Although the GP model attains the lowest error, the DMD variants are more computationally efficient because multi-step forecasts follow from spectral propagation (i.e., raising the eigenvalue matrix to the desired power). In contrast, a purely GP-based rollout must repeatedly lift each intermediate state, increasing computational overhead. Consequently, for long horizons, DMD methods often employ a reprojection scheme that combines both approaches; intermittent reprojections onto the state manifold mitigates drift arising from the fact that the learned finite-dimensional subspace may not be perfectly invariant under the true Koopman operator (Van Goor et al., 2023).
In Figure 4 we demonstrate how reprojections based on a mechanism defined on (2.12) can be effectively used to decrease the computational cost while also maintaining accuracy. Using the decoupled version of GP-DMD, the idea was simply to reproject when the Euclidean norm of the diagonal entries in ${\mathbf{K}}_{\textrm{pst}}^{k}\left({\bm{x}}_{0}\right)$ exceeded some predefined tolerance. At such points, the current estimate $\hat{\bm{x}}_{k}$ was treated as a noise-free measurement, and the feature vector $\mathbf{k}_{Z}(\hat{\bm{x}}_{k})$ was recomputed [^2] In evaluating ${\mathbf{\Xi}}_{\textrm{pst}}^{1}(\bm{x})={\kappa}_{\textrm{pst}}(\bm{x},\bm{x})\,\mathbf{K}_{\textrm{bc}}$ and subsequently (2.9), with $\sigma_{\mathcal{Y}}$ substituted for every $\sigma_{Y}$, we employed ${\mathbf{V}}_{\kappa}$ to bypass this costly computation..
@@ 3.4 Eigenfunction Uncertainty Quantification
Figures 5 and 6 show that the well locations align with regions of high certainty. This is because the dynamical systems spend most of their time around the wells, leading to dense sampling in these areas. While the kernel lengthscales also influence the geometry of the low-uncertainty regions, the dominant effect in these plots is the sampling distribution of the training data sets: initial conditions were chosen based on an optimized Latin hypercube design (Bates et al., 2004; Urquhart et al., 2020), and proceeded by rolling out trajectories with lengths of $20$ time-steps.
@@ 4 Conclusion
We have proposed a sparse kernel-based DMD algorithm which we formulated within a Bayesian framework that treats the embedded Perron–Frobenius operator (Klus et al., 2020b) as a random variable. The method extends EDMD by incorporating hyperparameter optimization, sparse dictionary learning, and enables propagating uncertainties in the eigenfunctions of the flow map.
The primary insight we have gained is that kernel transfer operators could benefit from a heteroskedastic observation model; one that is closely intertwined with a consistent notion of an operator-valued RKHS. However, we have another reason to modify our model with a more complex noise model; thus far, we have focused on compensating for measurement noise on the target variables. Yet, DMD is known to be susceptible to adverse effects arising from noise on the inputs (Dawson et al., 2016; Pan et al., 2021; Duke et al., 2012; Bagheri, 2014). Incidentally, one approach to dealing with input-noise is a heteroskedastic GP model (Lázaro-Gredilla and Titsias, 2011; McHutchon and Rasmussen, 2011; Wilson et al., 2012). Alternatively, in the Koopman community, several variants of DMD have been proposed to correct for the bias induced by input-noise (Hemati et al., 2017; Jiang and Liu, 2022; Scherl et al., 2020; Nonomura et al., 2019).
Earlier we posed the question of which philosophical stance is most appropriate for kernel transfer operators: a Bayesian or a frequentist perspective. We have argued that the Bayesian viewpoint offers ample opportunities to enrich the interpretation of a DMD models. At the same time, we have also demonstrated that diverse perspectives can be especially valuable, as uncovering the reasons for discrepancies between the two paradigms may lead to the insights that solve the problem. Thus, rather than exclusively favoring one approach above the other, we advocate for employing Bayesian and frequentist methods in tandem whenever possible.
@@ A.2 Bayes’ Law
The adjoint of the dictionary matrix is called the observation operator, since $\Phi_{X}^{\top}:\mathbb{H}\rightarrow\mathbb{R}^{N}$ is a mapping from the feature space to the observable space. By the reproducing property, the observation operator allows us to sample a function $f\in\mathbb{H}$. Therefore, we can define the vector of latent function values as $\bm{f}_{X}\coloneqq\Phi_{X}^{\top}f=[f(\bm{x}_{1}),\ldots,f(\bm{x}_{N})]^{\top}$, and the variance-covariance matrix by $\mathbf{K}_{XX}\coloneqq\big[\kappa_{\textrm{pr}}(\bm{x}_{i},\bm{x}_{j})\big]_{i,j=1}^{N}=\kappa_{\textrm{pr}}(\mathbf{X},\mathbf{X})=\Phi_{X}^{\top}\Phi_{X}$. Simply put, GPs treat every latent function value as a random variable.
@@
where ${\mathbf{A}}\coloneqq\tilde{\mathbf{K}}_{XX}^{-1}\mathbf{Y}^{\top}\in\mathbb{R}^{N}$ is the weight matrix, $\tilde{\mathbf{K}}_{XX}\coloneqq\mathbf{K}_{XX}+\sigma_{Y}^{2}\mathbf{I}_{N}$ is the covariance matrix for the noisy outputs [^3] The noise kernel models a constant covariance over the state space $\mathbb{X}$ with $\kappa_{\sigma}(\bm{x},\bm{x}^{\prime})\coloneqq\sigma^{2}\delta(\bm{x},\bm{x}^{\prime})$, which is a scaled Kronecker delta function. Couple this with the fact that the sum of kernel functions is a valid kernel., and $\mathbf{k}_{X}(\bm{x})\coloneqq\Phi_{X}^{\top}\varphi(\bm{x})=[\kappa_{\textrm{pr}}(\bm{x}_{1},\bm{x}),\ldots,\kappa_{\textrm{pr}}(\bm{x}_{N},\bm{x})]^{\top}$ is a feature vector. Now, all our questions about the noisy predictive posterior distribution at some unseen input $\bm{x}_{*}$ can be answered by querying the expression:
@@
Maximizing (A.4) naturally embodies Occam’s razor, and can be interpreted as engineering the prior GP such that it maximizes the volume of data-matching functions. [^4] One can often find an analytical expression for the gradient of (A.4) and use optimization algorithms such as conjugate gradient ascent or BFGS.
@@ A.3 Sparse Regression
Due to the Gaussian assumptions, the calculus of variations can be used to analytically solve for the optimal approximate posterior Gaussian process. This results in the posterior mean and covariance functions
$m_{\textrm{pst}}(\bm{x})$ $\coloneqq\mathbf{k}_{Z}^{\top}(\bm{x})\text{ }{\mathbf{A}},$ (A.5)
@@
Here, we are deviating from the notation in [Titsias, 2009] to highlight the similarities with (kernel) ridge-regression. The weight matrix, ${\mathbf{A}}\!\coloneqq\!\tilde{\mathbf{C}}_{XX}^{-1}\mathbf{K}_{ZX}\mathbf{Y}^{\top}$, contains the regression coefficients obtained from linear operations on matrices. The noisy Gramian is defined by $\tilde{\mathbf{C}}_{XX}\!\coloneqq\!\mathbf{K}_{ZX}\mathbf{K}_{ZX}^{\top}\!+\!\sigma_{Y}^{2}\mathbf{K}_{ZZ}$, and the cross-covariance matrix by $\mathbf{K}_{ZX}\!\coloneqq\!\kappa_{\textrm{pr}}(\mathbf{Z},\mathbf{X})\!\in\!\mathbb{R}^{M\times N}$. Lastly, to compute the posterior kernel the weights for the information gain term are $\tilde{\mathbf{B}}\!\coloneqq\!\left(\mathbf{K}_{ZZ}^{-1}-\sigma_{Y}^{2}\!\tilde{\mathbf{C}}_{XX}^{-1}\right)$.
@@ B.1 The Koopman and Perron–Frobenius Operators
The starting point for Koopman theory is mapping the RDS from the lower-dimensional nonlinear space $\mathbb{X}$ to a higher-dimensional linear space $\mathbb{G}$, such that no information is lost when moving to this feature space. In short, operator theory can be viewed as a trade-off between lifting the state space into a feature space with more complex states but simpler dynamics. To perform the lifting, the literature defines an observable function as, $g\in\mathbb{G}$, where $\mathbb{G}=\mathcal{L}^{\infty}(\mathbb{X})$. [^6] Other Banach spaces for $\mathbb{G}$ are also valid, e.g., the Hilbert space of Lebesgue square-integrable functions, i.e., $\mathbb{G}\coloneqq L^{2}(\mathbb{X})$ [Klus et al., 2020b, Ikeda et al., 2022].
@@
Hence we call any operator, transporting some object (a distribution, or an observable) by the dynamics, a transfer operator, i.e., the action of the process on functions of the state. The mathematical formulations of kernel transfer operators follow from the assumptions that the observables are elements of an RKHS feature space, and the operators are compact with a discrete spectrum [Klus et al., 2020b]. This implies that we may set $\mathbb{G}=\mathbb{H}$.
@@ Appendix C Embedding Probability Distributions
By construction, KME allows for featurization of the output variable, while GP regression is typically presented with scalar-valued outputs [^7] In this work, we of course show, how this is not true in general.. In other words, the output space $\mathbb{Y}$ is not limited to $\mathbb{Y}\subseteq\mathbb{R}$ [Song et al., 2009]. In contrast with (A.2) we want to model $\mathbb{E}_{Y\mid\bm{x}}[g(Y)\mid X=\bm{x}],\;g\in\mathbb{H}_{Y}$. Apart from this generalization, the mathematical formalisms for the predictive mean (A.2) and the kernel conditional mean embedding (C.2) are exactly the same, as are the predictive variance (A.3) and the Mahalanobis distance [Chowdhury and Gopalan, 2017, Chowdhury and Oliveira, 2023]. It is a case where the GP regression and KME communities, have developed overlapping tools while having different goals in mind.
A limitation of the KME approach is that it recovers only part of the structure available in GP regression: in particular, concepts such as the marginal likelihood do not translate as directly into the embedding framework.
@@ C.2 Joint Distributions
Therefore, $\mathcal{C}_{XX}$ is a self-adjoint operator and $\mathcal{C}_{YX}$ is the adjoint of $\mathcal{C}_{XY}$. Note that the joint feature map is a rank one operator from $\mathbb{H}_{X}$ to $\mathbb{H}_{Y}$. Consequently, there is an interesting duality of perspectives at play, because just like the mean embedding was an element in the RKHS $\mu_{X}\in\mathbb{H}_{X}$, so the cross-covariance operator is an element in the tensor product feature space $\mathcal{C}_{YX}\in\mathbb{H}_{YX}$.
@@ C.3 Conditional Distributions
The condition $\mathbb{E}_{Y\mid X}\left[g(Y)\mid X=\cdot\right]\in\mathbb{H}_{X}\;\forall\;g\in\mathbb{H}_{Y}$, will always hold for finite domains with characteristic kernels, but may not be valid for a continuous domain [Fukumizu et al., 2013]. A common approach to extend the condition to other systems is to consider the regularized inverse $(\mathcal{C}_{XX}+\varepsilon\mathcal{I})^{-1}$, where $\mathcal{I}$ is the identity operator on $\mathbb{H}_{X}$, and $\varepsilon>0$ is a Tikhonov regularization parameter [Tikhonov, 1977]. Moreover, $\mathcal{C}_{XX}$ is a compact trace class operator, with eigenvalues that accumulate at zero when $\mathbb{H}_{X}$ is infinite dimensional [Hsing and Eubank, 2015]. With the regularization, the inverse becomes well-posed.

### v2

@@ 3.2 Decoupling the Lifted Observation Model
To assess the sensitivity of our algorithm to the approximation introduced in Definition 2.1, we separately train a regularization parameter for the lifted model using the kernelized targets $\mathbf{K}_{ZY}$ (after estimating the optimal kernel hyperparameters and pseudo-inputs as discussed in Section 2.5). In effect, we are defining the RKHS noise model with $\sigma_{\mathcal{Y}}^{2}\kappa_{\textrm{bc}}\left(\cdot,\cdot\right)$ where $\sigma_{\mathcal{Y}}^{2}$ is not directly coupled to the variance of the sensor noise in the state space, $\sigma_{Y}^{2}$. Put differently, the regularization parameter in (2.5) now corresponds to $\sigma_{\mathcal{Y}}$, whereas in (2.7) it is still $\sigma_{Y}$. This establishes a division of roles between the two regularization terms: $\sigma_{Y}$ is predominantly responsible for projecting to the state space, while $\sigma_{\mathcal{Y}}$ controls the accuracies of the multi-step forecasts. Consequently, the relationship in (2.11) does not hold exactly anymore, but our model is more flexible.
The effect of decoupling is illustrated in Figure 3 where we observe that the growth rate of the GP-TCCA model has markedly reduced, leading to more accurate long-term predictions. We therefore increase the credence in our hypothesis: rigorously accounting for the distortion introduced by the nonlinear feature map will yield a more robust model. For example, we are naturally tempted to model the statistical properties of $\nu_{i}$ as non-Gaussian and input-dependent (i.e., heteroskedastic). This is apparent when we notice that $\kappa_{\mathrm{pr}}(\bm{x},Y)$ is typically a nonlinear many-to-one transformation of $Y$.
@@ 3.3 Reprojections
Although the GP model attains the lowest error, the DMD variants are more computationally efficient, because multi-step forecasts follow from spectral propagation (i.e., raising the eigenvalue matrix to the desired power). In contrast, a purely GP-based rollout must repeatedly lift each intermediate state, increasing computational overhead. Consequently, for long horizons, DMD methods often employ a reprojection scheme that combines both approaches; intermittent reprojections onto the state manifold mitigates drift arising from the fact that the learned finite-dimensional subspace may not be perfectly invariant under the true Koopman operator (Van Goor et al., 2023).
In Figure 4 we demonstrate how reprojections based on a mechanism defined on (2.12) can be effectively used to decrease the computational cost while also maintaining accuracy. The idea was simply to reproject when the Euclidean norm of the diagonal entries in ${\mathbf{K}}_{\textrm{pst}}^{k}\left({\bm{x}}_{0}\right)$ exceeded some predefined tolerance. At such points, the current estimate $\hat{\bm{x}}_{k}$ was treated as a noise-free measurement, and the feature vector $\mathbf{k}_{Z}(\hat{\bm{x}}_{k})$ was recomputed [^5] In evaluating ${\mathbf{\Xi}}_{\textrm{pst}}^{1}(\bm{x})={\kappa}_{\textrm{pst}}(\bm{x},\bm{x})\,\mathbf{K}_{\textrm{bc}}$ and subsequently (2.9), we employed spectral propagation with ${\mathbf{V}}_{\kappa}$ to compute the feature vector $\mathbf{k}_{Z}(\hat{\bm{x}}_{k})$, thereby bypassing multiple evaluations of a potentially costly kernel function..
@@ 3.4 Eigenfunction Uncertainty Quantification
Figures 5 and 6 show that the well locations align with regions of high certainty. This is because the dynamical systems spend most of their time around the wells, leading to dense sampling in these areas. While the kernel lengthscales also influence the geometry of the low-uncertainty regions, the dominant effect in these plots is the sampling distribution of the training data sets: initial conditions were chosen based on an optimized Latin hypercube design (Bates et al., 2004; Urquhart et al., 2020), and succeeded by rolling out trajectories with lengths of $20$ time-steps.
@@ 4 Conclusion
We have proposed a sparse kernel-based DMD algorithm which we formulated within a Bayesian framework that treats the embedded Perron–Frobenius operator (Klus et al., 2020c) as a random variable. By unifying operator learning with the mature field of Gaussian process regression the method extends EDMD by incorporating hyperparameter optimization, sparse dictionary learning, enables propagating uncertainties in the eigenfunctions of the flow map, and provides a criteria for reprojections.
The primary insight we have gained from the ablation studies is that kernel transfer operators could benefit from a non-Gaussian and a heteroskedastic observation model; one that is closely intertwined with a Bayesian consistent notion of an operator-valued RKHS. However, we have another reason to modify our model with a more complex noise model; thus far, we have only focused on compensating for measurement noise on the target variables. Yet, DMD is known to be susceptible to adverse effects arising from noise on the inputs (Dawson et al., 2016; Pan et al., 2021; Duke et al., 2012; Bagheri, 2014). Incidentally, one approach to dealing with input-noise is a heteroskedastic GP model (Lázaro-Gredilla and Titsias, 2011; McHutchon and Rasmussen, 2011; Wilson et al., 2012). Alternatively, in the Koopman community, several variants of DMD have been proposed to correct for the bias induced by input-noise (Hemati et al., 2017; Jiang and Liu, 2022; Scherl et al., 2020; Nonomura et al., 2019).
Earlier we posed the question of which philosophical stance is most appropriate for kernel transfer operators: a Bayesian or a frequentist perspective. We have argued that the Bayesian viewpoint offers ample opportunities to enrich the interpretation of DMD models. At the same time, we have also demonstrated that diverse perspectives can be especially valuable, as uncovering the reasons for discrepancies between the two paradigms led to the insights that solved the problem. Thus, rather than exclusively favoring one approach above the other, we advocate for employing Bayesian and frequentist methods in tandem whenever possible.
@@ A.2 Bayes’ Law
The adjoint of the dictionary operator is called the observation operator, since $\Phi_{X}^{\top}:\mathbb{H}\rightarrow\mathbb{R}^{N}$ is a mapping from the feature space to the observable space. By the reproducing property, the observation operator allows us to sample a function $f\in\mathbb{H}$. Therefore, we can define the vector of latent function values as $\bm{f}_{X}\coloneqq\Phi_{X}^{\top}f=[f(\bm{x}_{1}),\ldots,f(\bm{x}_{N})]^{\top}$, and the variance-covariance matrix by $\mathbf{K}_{XX}\coloneqq\big[\kappa_{\textrm{pr}}(\bm{x}_{i},\bm{x}_{j})\big]_{i,j=1}^{N}=\kappa_{\textrm{pr}}(\mathbf{X},\mathbf{X})=\Phi_{X}^{\top}\Phi_{X}$. Simply put, GPs treat every latent function value as a random variable.
@@
where ${\mathbf{A}}\coloneqq\mathbf{Y}\>\tilde{\mathbf{K}}_{XX}^{-1}\in\mathbb{R}^{1\times N}$ is the weight matrix, $\tilde{\mathbf{K}}_{XX}\coloneqq\mathbf{K}_{XX}+\sigma_{Y}^{2}\mathbf{I}_{N}$ is the covariance matrix for the noisy outputs [^6] The noise kernel models a constant covariance over the state space $\mathbb{X}$ with $\kappa_{\sigma}(\bm{x},\bm{x}^{\prime})\coloneqq\sigma^{2}\delta(\bm{x},\bm{x}^{\prime})$, which is a scaled Kronecker delta function. Couple this with the fact that the sum of kernel functions is a valid kernel., and $\mathbf{k}_{X}(\bm{x})\coloneqq\Phi_{X}^{\top}\varphi(\bm{x})=[\kappa_{\textrm{pr}}(\bm{x}_{1},\bm{x}),\ldots,\kappa_{\textrm{pr}}(\bm{x}_{N},\bm{x})]^{\top}$ is a feature vector. Now, all our questions about the noisy predictive posterior distribution at some unseen input $\bm{x}_{*}$ can be answered by querying the expression:
@@
Maximizing (A.4) naturally embodies Occam’s razor, and can be interpreted as engineering the prior GP such that it maximizes the volume of data-matching functions. [^7] One can often find an analytical expression for the gradient of (A.4) and use optimization algorithms such as conjugate gradient ascent or BFGS [Zhu et al., 1997].
@@ A.3 Sparse Regression
Due to the Gaussian assumptions, the calculus of variations can be used to analytically solve for the optimal approximate posterior Gaussian process. This results in the following posterior mean and covariance functions:
$m_{\textrm{pst}}(\bm{x})$ $\coloneqq{\mathbf{A}}\>\mathbf{k}_{Z}(\bm{x}),$ (A.5)
@@
Here, we are deviating from the notation in [Titsias, 2009] to highlight the similarities with (kernel) ridge-regression. The weight matrix, ${\mathbf{A}}\!\coloneqq\!\mathbf{Y}\,\mathbf{K}_{ZX}^{\top}\,\tilde{\mathbf{C}}_{XX}^{-1}$, contains the regression coefficients obtained from linear operations on matrices. The noisy Gramian is defined by $\tilde{\mathbf{C}}_{XX}\!\coloneqq\!\mathbf{K}_{ZX}\mathbf{K}_{ZX}^{\top}\!+\!\sigma_{Y}^{2}\mathbf{K}_{ZZ}$, and the cross-covariance matrix by $\mathbf{K}_{ZX}\!\coloneqq\!\kappa_{\textrm{pr}}(\mathbf{Z},\mathbf{X})\!\in\!\mathbb{R}^{M\times N}$. Lastly, to compute the posterior kernel the weights for the information gain term are $\tilde{\mathbf{B}}\!\coloneqq\!\left(\mathbf{K}_{ZZ}^{-1}-\sigma_{Y}^{2}\!\tilde{\mathbf{C}}_{XX}^{-1}\right)$.
@@ B.1 The Koopman and Perron–Frobenius Operators
The starting point for Koopman theory is mapping the RDS from the lower-dimensional nonlinear space $\mathbb{X}$ to a higher-dimensional linear space $\mathbb{G}$, such that no information is lost when moving to this feature space. In short, operator theory can be viewed as a trade-off between lifting the state space into a feature space with more complex states but simpler dynamics. To perform the lifting, the literature defines an observable function as, $g\in\mathbb{G}$, where $\mathbb{G}=\mathcal{L}^{\infty}(\mathbb{X})$. [^9] Other Banach spaces for $\mathbb{G}$ are also valid, e.g., the Hilbert space of Lebesgue square-integrable functions, i.e., $\mathbb{G}\coloneqq L^{2}(\mathbb{X})$ [Klus et al., 2020c, Ikeda et al., 2022].
@@
Hence we call any operator, transporting some object (a distribution, or an observable) by the dynamics, a transfer operator, i.e., the action of the process on functions of the state. The mathematical formulations of kernel transfer operators follow from the assumptions that the observables are elements of an RKHS feature space, and the operators are compact with a discrete spectrum [Klus et al., 2020c]. This implies that we may set $\mathbb{G}=\mathbb{H}$.
@@ Appendix C Embedding Probability Distributions
From the outset, KME allows for featurization of the output variable, while GP regression is typically presented with pure scalar-valued outputs [^10] In this work, we bridge the gap by providing a Bayesian interpretation of $\varphi(Y)$.. In other words, the output space $\mathbb{Y}$ is not limited to $\mathbb{Y}\subseteq\mathbb{R}$ [Song et al., 2009]. In contrast with (A.2) we directly model $\mathbb{E}_{Y\mid\bm{x}}[g(Y)\mid X=\bm{x}],\;g\in\mathbb{H}_{Y}$. Apart from this, the mathematical formalisms for the predictive mean (A.2) and the kernel conditional mean embedding (C.2) are exactly the same, as are the predictive variance (A.3) and the Mahalanobis distance [Chowdhury and Gopalan, 2017, Chowdhury and Oliveira, 2023]. It is a case where the GP regression and KME communities, have developed overlapping tools while having different goals in mind.
A limitation of the KME approach is that it recovers only part of the structure available in GP regression: in particular, other concepts such as the marginal likelihood do not translate directly into the embedding framework.
@@ C.2 Joint Distributions
Therefore, $\mathcal{C}_{XX}$ is a self-adjoint operator and $\mathcal{C}_{YX}$ is the adjoint of $\mathcal{C}_{XY}$. Note that the joint feature map is a rank one operator. Consequently, there is an interesting duality of perspectives at play, because just like the mean embedding was an element in the RKHS $\mu_{X}\in\mathbb{H}_{X}$, so the cross-covariance operator is an element in the tensor product feature space $\mathcal{C}_{YX}\in\mathbb{H}_{YX}$.
@@ C.3 Conditional Distributions
The condition $\mathbb{E}_{Y\mid X}\left[g(Y)\mid X=\cdot\right]\in\mathbb{H}_{X}\;\forall\;g\in\mathbb{H}_{Y}$, will always hold for finite domains with characteristic kernels, but may not be valid for a continuous domain [Fukumizu et al., 2013]. A common approach to extend the condition to other systems is to consider the regularized inverse $(\mathcal{C}_{XX}+\varepsilon\mathcal{I})^{-1}$, where $\mathcal{I}$ is the identity operator on $\mathbb{H}_{X}$, and $\varepsilon>0$ is a Tikhonov regularization parameter [Tikhonov, 1977]. Moreover, $\mathcal{C}_{XX}$ is a compact trace class operator, with eigenvalues that accumulate at zero when $\mathbb{H}_{X}$ is infinite dimensional [Hsing and Eubank, 2015]. The regularization ensures a well-posed inverse.